OpenStudy (anonymous):

differentiate y= (e^-2x) ln(x^2)

OpenStudy (anonymous):

is it \[y=e^{-2x}\ln(x^2)\]?

OpenStudy (anonymous):

yupp haha

zepdrix (zepdrix):

Sooo product rule, yes? :) \[\Large y'\quad=\quad \color{royalblue}{\left[e^{-2x}\right]'}\ln(x^2)+e^{-2x}\color{royalblue}{\left[\ln(x^2)\right]'}\]

OpenStudy (anonymous):

ohh okay that clears things up a bit:) haha but I don't know how to find the first derivatives of the two

zepdrix (zepdrix):

Hmm, remember this derivative?\[\Large (e^x)'\quad=\quad ?\]

OpenStudy (anonymous):

e^x so does e^-2x = -2e^-2x and ln(x^2) become 1/(x^2)?

zepdrix (zepdrix):

Hmm your first one looks good.\[\Large y'\quad=\quad \color{orangered}{\left[-2e^{-2x}\right]}\ln(x^2)+e^{-2x}\color{royalblue}{\left[\ln(x^2)\right]'}\]

zepdrix (zepdrix):

\[\Large \left[\ln(x^2)\right]'\quad=\quad \frac{1}{x^2}(x^2)' \quad=\quad \frac{2}{x}\]

zepdrix (zepdrix):

You could also use rules of exponents to do that if you needed :)\[\Large \left[\ln(x^2)\right]'\quad=\quad\left[2\ln(x)\right]'\quad=\quad2\left[\ln(x)\right]'\quad=\quad 2\frac{1}{x}\]